Determine the pH of a 0.22 M NaF solution. The Ka of HF is 3.5 × 10-5
Answer #1

have : [NaF] = 0.22 M

Ka of HF is 3.5 × 10-5

pH =?

from:

NaF(s) dissolve in water give: Na+(aq)+F-(aq)

thus,F-(aq)+H2O(l) <---------->HF(aq)+OH-(aq)

0.22 M00

if0.22 - BBB(mol/L)

Ka = [HF].[OH-]/[F-], but [OH-] is less we give [F-] = 0.22 - B≈ 0.22 M

3.5 x 10-5 = B2/0.22 --> B = [OH-] = (3.5 x 10-5x0.22)1/2≈2.775x10-3 M

we have; pOH = -log[OH-] and pH + pOH = 14

-> pH = 14 -pOH = 14 + log[OH-] = 14 - log(2.775x10-3)≈ 11.443

thus, the pH of a 0.22 M NaF is pH = 11.443